### v1

@@ 3.4 A quasi-best approximation result for the Scott-Zhang operator
Let $\Phi$ be a given $N$-function that satisfies the conditions in Assumption 2.1.4, let $\omega\in A_{1}$, and let $v\in W^{1,\Phi}(\omega,\Omega)$. Then, for each $T\in\mathscr{T}_{h}$ we have the bound
@@
where the hidden constant depends on $\sigma$, $\Delta_{2}(\Phi)$, and $[\omega]_{A_{1}}$. In particular, if $\nabla\mathbf{V}(\nabla v)$ belongs to $L^{2}(\omega,\Omega)$, then we have the bound
@@
where the hidden constant depends on $\sigma$, $\Delta_{2}(\Phi)$, and $[\omega]_{A_{1}}$.
Proof 3.8.
@@
We now apply the stability estimate of Corollary 33 with $a=|\mathbf{q}|$ to the function $\mathfrak{p}-v$ and arrive at
@@
In what follows, we bound $\mathbf{I}_{T}$. To do this, we use that $T\subset S_{T}$ with $|T|\simeq|S_{T}|$ to obtain
@@
A combination of the bounds derived for $\mathbf{I}_{T}$ and $\mathbf{II}_{T}$ allows us to conclude for every $\mathbf{q}\in\mathbb{R}^{d}$ that
@@ 3.5 An a priori error bound
Theorem 3.9 (a priori error bound).
Let $u\in W_{0}^{1,\varphi}(\Omega)$ be the weak solution of (27), and let $u_{h}\in\mathbb{V}_{h}$ be its finite element approximation defined in (28), where $\varphi$ satisfies the conditions in Assumption 2.1.4. If $\nabla\mathbf{V}(\nabla u)\in L^{2}(\omega,\Omega)$, then
@@
where the hidden constant depends on $\sigma$, $\Delta_{2}(\Phi)$, and $[\omega]_{A_{1}}$. Recall that $\omega\in A_{1}$ and that $\varphi$ is an $N$-function that satisfies the conditions in Assumption 2.1.4.
Proof 3.10.
The error bound (37) follows from the quasi-best approximation property (30) in conjunction with the the error estimate (35).
@@
Note that it is fundamental to the error estimate of Theorem 3.9 that $\nabla{\mathbf{V}}(\nabla u)\in L^{2}(\omega,\Omega)$. In the unweighted setting, i.e., $\omega\equiv 1$, such results are proved in [1, 7, 8, 15, 18].
@@ 4 The obstacle problem
We now consider the constrained minimization of the functional $\mathcal{J}$ defined in (1). In the same setting and with the same assumptions as in Section 3, we additionally assume that we are given $\psi\in C^{2}(\bar{\Omega})$ with $\psi\leq 0$ on $\partial\Omega$. We define the admissible set
[…]
Recall that $\omega\in A_{1}$ and that $\varphi$ is an $N$-function that satisfies the conditions in Assumption 2.1.4. The problem that interests us is to find $u\in\mathcal{K}_{\psi}$ such that
$\mathcal{J}(u)\leq\mathcal{J}(v)\qquad\forall v\in\mathcal{K}_{\psi}.$
A standard convex minimization argument (see also [21, 27]) shows that the solution to this problem can be equivalently characterized as the following variational inequality. Find $u\in\mathcal{K}_{\psi}$ such that
@@
Since it will be useful for what follows, we define $\lambda\in\left(W^{1,\varphi}(\omega,\Omega)\right)^{\prime}$ as
@@
Note, then, that (39) can be rewritten as follows: Find $u\in\mathcal{K}_{\psi}$ such that
@@
  • For every $\phi\in C_{0}^{\infty}(\Omega)$ that is nonnegative, i.e., $\phi\geq 0$ in $\Omega$, we have
    $\langle\lambda,\phi\rangle\geq 0.$
  • The following complementarity condition holds:
    $\langle\lambda,u-\psi\rangle=0.$
Proof 4.2.
Let $\phi\in C_{0}^{\infty}(\Omega)$ with $\phi\geq 0$ in $\Omega$. Define $v=u+\phi\in W^{1,\varphi}_{0}(\omega,\Omega)$. Since $v$ satisfies $v\geq u\geq\psi$ in $\Omega$, $v$ is admissible in (39). We can therefore substitute $v$ in the variational inequality (39) and obtain
$0\geq\langle\lambda,u-(u+\phi)\rangle=-\langle\lambda,\phi\rangle.$
This proves the first property. We also note that this property shows that $\lambda$ defines a nonnegative distribution. The Riesz-Schwartz theorem (see [41, Chapter 1, §4, Théorème V] and [44, Chapter 1, Sec. 1.7, Theorem II]) then shows that $\lambda$ is a nonnegative Radon measure.
@@
Now let $\mathcal{N}\subset\Omega^{+}$ be open, and let $\phi\in C_{0}^{\infty}(\mathcal{N})$. For $\epsilon>0$, but sufficiently small, we have that the function $v=u-\epsilon\phi\in W^{1,\varphi}_{0}(\omega,\Omega)$ belongs to $\mathcal{K}_{\psi}$. In fact, we have that $u(x)>\psi(x)$ up to a null set on $\Omega^{+}$. Then, for a sufficiently small $\epsilon>0$, we obtain that $u(x)-\epsilon\phi(x)\geq\psi(x)$ in $\Omega$. If we replace this function $v$ in (39), we obtain
@@
$\int_{\Omega^{+}}\phi(x)\,\mbox{\rm d}\lambda(x)=0,\qquad\forall\phi\in C_{0}^{\infty}(\Omega^{+}):\ \phi\geq 0\text{ in }\Omega^{+}.$
@@ 4.1 Finite element discretization
To approximate the solution of (39), we retain the notation and constructions from Section 3.1. In addition, we need to introduce the positivity preserving interpolant $\wp_{h}:L^{1}(\Omega)\to\mathbb{V}_{h}$, which was originally developed in [12]. Let $\mathscr{V}=\{{\mathtt{v}}\}$ be the collection of vertices of $\mathscr{T}_{h}$ and let $\mathscr{V}_{\mathrm{in}}\coloneqq\mathscr{V}\cap\Omega$. Recall that since $\mathbb{V}_{h}$ consists of piecewise linears, there is a bijection between $\mathscr{V}_{\mathrm{in}}$ and the canonical Lagrange basis $\{\phi_{\mathtt{v}}\}_{{\mathtt{v}}\in\mathscr{V}_{\mathrm{in}}}$ of $\mathbb{V}_{h}$. Given ${\mathtt{v}}\in\mathscr{V}_{\mathrm{in}}$, we let
@@
Standard arguments of variational inequalities yield the existence and uniqueness of a solution to (44).
@@ 4.2 A priori error bounds
Let $u\in\mathcal{K}_{\psi}$ and $u_{h}\in\mathcal{K}_{h,\psi}$ be the solutions of (39) and (44), respectively. If $u\in L^{2}(\omega,\Omega)$ and $\lambda\in L^{2}(\omega^{-1},\Omega)$, then we have
@@
Proof 4.4.
We begin the proof by setting some notation. We define the error $e=u-u_{h}$ and the auxiliary function $z=u-\wp_{h}u$. Note that $\wp_{h}u\in\mathcal{K}_{h,\psi}$.
@@
$\|{\mathbf{V}}(\nabla u)-{\mathbf{V}}(\nabla u_{h})\|_{L^{2}(\omega,\Omega)}^{2}\lesssim\int_{\Omega}\omega(x)\left[{\mathbf{A}}(\nabla u(x))-{\mathbf{A}}(\nabla u_{h}(x))\right]\cdot\nabla e(x)\,\mbox{\rm d}x\eqqcolon\mathfrak{D},$
which we further decompose as
@@
    $\leq\|\lambda\|_{L^{2}(\omega^{-1},\Omega)}\|\left[u-\psi\right]-\wp_{h}\left[u-\psi\right]\|_{L^{2}(\omega,\Omega)},$
@@
$\mathfrak{L}\leq C\delta\|{\mathbf{V}}(\nabla u)-{\mathbf{V}}(\nabla u_{h})\|_{L^{2}(\omega,\Omega)}^{2}+C_{\delta}\int_{\Omega}\omega(x)\varphi_{|\nabla u(x)|}(|\nabla z(x)|)\,\mbox{\rm d}x.$
If we substitute this estimate and the one obtained for $\mathfrak{N}$ in the bound in which $\mathfrak{D}$ is defined, we obtain
(45) $\|{\mathbf{V}}(\nabla u)-{\mathbf{V}}(\nabla u_{h})\|_{L^{2}(\omega,\Omega)}^{2}\leq C\delta\|{\mathbf{V}}(\nabla u)-{\mathbf{V}}(\nabla u_{h})\|_{L^{2}(\omega,\Omega)}^{2}\\ +C_{\delta}\int_{\Omega}\omega(x)\varphi_{|\nabla u(x)|}(|\nabla z(x)|)\,\mbox{\rm d}x+C\|\lambda\|_{L^{2}(\omega^{-1},\Omega)}\|\left[u-\psi\right]-\wp_{h}\left[u-\psi\right]\|_{L^{2}(\omega,\Omega)}.$
The term $\|{\mathbf{V}}(\nabla u)-{\mathbf{V}}(\nabla u_{h})\|_{L^{2}(\omega,\Omega)}^{2}$ appearing in the previous inequality can be absorbed on the left hand side if $\delta$ is chosen carefully. This shows that
(46) $\|{\mathbf{V}}(\nabla u)-{\mathbf{V}}(\nabla u_{h})\|_{L^{2}(\omega,\Omega)}^{2}\lesssim\|\lambda\|_{L^{2}(\omega^{-1},\Omega)}\|\left[u-\chi\right]-\wp_{h}\left[u-\chi\right]\|_{L^{2}(\omega,\Omega)}\\ +\int_{\Omega}\omega(x)\varphi_{|\nabla u(x)|}(|\nabla z(x)|)\,\mbox{\rm d}x,$
with $z=u-\wp_{h}u$. Use the equivalences (15)–(16) one last time to obtain the assertion. This concludes the proof.
@@ 4.3 The positivity preserving interpolant
Note that up to this point, no property other than positivity has been required of $\wp_{h}$. In what follows, however, the stability and approximation properties of $\wp_{h}$ are required. Ideally, we would like to have an operator that is a projection and satisfies (31). Note, however, that $\wp_{h}$ is not a projection. In fact, it is not possible to construct a positivity preserving projection; see [38]. For this reason, we need to prove suitable properties for $\wp_{h}$.
@@
    $\wp_{h}$ is weighted Orlicz stable: Let $\omega\in A_{1}$ and let $\Phi$ be an $N$-function that satisfies the conditions in Assumption 2.1.4. Then, for every $T\in\mathscr{T}_{h}$ and for all $a\geq 0$,
@@
    $\wp_{h}$ has weighted approximation properties: If $w\in W^{1,2}_{0}(\omega,\Omega)$, then, for all $T\in\mathscr{T}_{h}$,
@@
Proof 4.6.
We examine each statement separately.
@@
    Since $w_{|\mathcal{B}_{\mathtt{v}}}\in\mathbb{P}_{1}$, we can write $w(x)=a+\mathbf{b}\cdot(x-{\mathtt{v}})$ for some $a\in\mathbb{R}$ and $\mathbf{b}\in\mathbb{R}^{d}$. Then, we have
@@
    This follows from repeating the previous computation on every vertex of $T$, which forms an unisolvent set for $\mathbb{P}_{1}(T)$.
@@
    A suitable choice of $\mathfrak{p}$ then shall imply the thesis.
    To prove (47), we proceed differently depending on whether $T$ is an interior element or $T$ touches the boundary. If $T$ is an interior element, i.e., $T\cap\partial\Omega=\emptyset$, then local invariance and stability imply the result. On the other hand, if the element $T$ touches the boundary we proceed as follows. Let $\{{\mathtt{v}}_{i}^{T}\}_{i=0}^{d}$ be the collection of vertices of $T$. Since $(\wp_{h}w-\mathfrak{p})_{|T}\in\mathbb{P}_{1}$, we can write
@@
    We now use that $\|\phi_{\mathtt{v}}\|_{L^{\infty}(\Omega)}=1$ for all ${\mathtt{v}}\in\mathscr{V}_{\mathrm{in}}$ to obtain the estimate
@@
    On the other hand, if ${\mathtt{v}}_{i}^{T}\in\partial\Omega$, then $\wp_{h}w({\mathtt{v}}_{i}^{T})=0$ by definition. Let $F\subset T$ be the $(d-1)$-dimensional subsimplex containing ${\mathtt{v}}_{i}^{T}$. We thus have that
@@
Proposition 4.7 (interpolation error).
Let $w\in W^{1,\varphi}_{0}(\omega,\Omega)$ verify $\nabla{\mathbf{V}}(\nabla w)\in L^{2}(\omega,\Omega)$. Assume that $\mathscr{T}_{h}=\{T\}$ is such that no simplex $T$ has more than one $(d-1)$-dimensional subsimplex on $\partial\Omega$ and that $\omega\in A_{1}(\Omega)$. If $h$ is sufficiently small, then
@@
where the hidden constant depends on $[\omega]_{A_{1}}$ and $\|\omega\|_{L^{\infty}({\mathcal{G}})}$ but it is independent of $h$. Recall that $\varphi$ is an $N$-function that satisfies the conditions in Assumption 2.1.4.
Proof 4.8.
@@
Let $T\in\mathscr{T}_{h}^{\mathrm{in}}$. The local invariance of $\wp_{h}$, proved in Lemma 4.5, implies $\wp_{h}\mathfrak{p}=\mathfrak{p}$. Using the equivalences (15)–(16) and the weighted Orlicz stability of $\wp_{h}$ derived in Lemma 4.5, we obtain the estimate
$\mathbf{I}\mathbf{I}$ $\lesssim\int_{T}\omega(x)\varphi_{|\nabla\mathfrak{p}|}(|\nabla\wp_{h}(w-\mathfrak{p})|)\,\mbox{\rm d}x\lesssim\int_{S_{T}}\omega(x)\varphi_{|\nabla\mathfrak{p}|}(|\nabla(w-\mathfrak{p})(x)|\,\mbox{\rm d}x$
@@
As a result, for each $T\in\mathscr{T}_{h}^{\mathrm{in}}$ we have the bound
@@
where $\mathfrak{p}\in\mathbb{P}_{1}(\mathbb{R}^{d})$ is arbitrary. Let $\mathbf{Q}\in\mathbb{R}^{d}$ and choose $\mathbf{z}=\nabla\mathfrak{p}\in\mathbb{R}^{d}$ so that ${\mathbf{V}}(\mathbf{z})=\mathbf{Q}$. This is possible because ${\mathbf{V}}$ is surjective. This gives us, for every $T\in\mathscr{T}_{h}^{\mathrm{in}}$,
@@
We now consider $T\in\mathscr{T}_{h}^{\partial}$. If $h$ is sufficiently small, there is no fault in assuming that $T\subset{\mathcal{G}}$. Recall that ${\mathcal{G}}$ is as in Definition 2.5. Let $F\subset T$ be the unique $(d-1)$-dimensional subsimplex such that $F=T\cap\partial\Omega$ and denote by $\boldsymbol{\zeta}$ the unit outer normal to $T$ on $F$. Now let $\{{\mathtt{v}}_{i}^{T}\}_{i=0}^{d}$ be the set of vertices of $T$ and assume that $\{{\mathtt{v}}_{i}^{T}\}_{i=0}^{d}$ is numbered such that $F=\operatorname{conv}\{{\mathtt{v}}_{i}^{T}\}_{i=0}^{d-1}$. Define $\mathfrak{p}\in\mathbb{P}_{1}(\mathbb{R}^{d})$ as $\mathfrak{p}(x)=b\boldsymbol{\zeta}\cdot(x-{\mathtt{v}}_{0}^{T})$, where $b\in\mathbb{R}$ is to be chosen. We note that, by construction, $\mathfrak{p}_{|F}=0$. Moreover, we have by definition that $\wp_{h}\mathfrak{p}_{|F}=0$. Finally, the symmetry of $\wp_{h}$ proved in Lemma 4.5 implies that $\wp_{h}\mathfrak{p}({\mathtt{v}}_{d}^{T})=\mathfrak{p}({\mathtt{v}}_{d}^{T})$. From this follows that $\wp_{h}\mathfrak{p}_{|T}=\mathfrak{p}$.
Based on the constructions described in the previous paragraph, the weighted Orlicz stability of $\wp_{h}$ (see Lemma 4.5) allows us to treat $\mathbf{I}\mathbf{I}$ as before and obtain that, for every $T\in\mathscr{T}_{h}^{\partial}$,

### v2

@@ 4.3 A quasi-best approximation result for the Scott-Zhang operator
Let $\Phi$ be an $N$-function satisfying Assumption 3, let $\omega\in A_{\Phi}$, and let $v\in W^{1,\Phi}(\omega,\Omega)$. Then, for each $T\in\mathscr{T}_{h}$, we have
@@
where the hidden constant depends on $\sigma$, $\omega]_{A_{i(\Phi)}}$, and the characteristics of $\Phi$. In particular, if $\nabla\mathbf{V}(\nabla v)$ belongs to $L^{2}(\omega,\Omega)$, then we have the bound
@@
where the hidden constant depends on $\sigma$, $\omega]_{A_{i(\Phi)}}$, and the characteristics of $\Phi$.
Proof 4.8.
@@
We now apply Corollary 40 with $a=|\mathbf{q}|$ to the function $\mathfrak{p}-v$ and arrive at
@@
We now bound $\mathbf{I}_{T}$. To do this, we use that $T\subset S_{T}$ with $|T|\simeq|S_{T}|$ to obtain
@@
Combining the bounds for $\mathbf{I}_{T}$ and $\mathbf{II}_{T}$ we conclude that, for every $\mathbf{q}\in\mathbb{R}^{d}$,
@@ 4.4 An error estimate
Theorem 4.9 (error estimate).
Let $\Phi$ be an $N$-function satisfying Assumption 3, and let $\omega\in A_{\Phi}$. Let $u\in W_{0}^{1,\Phi}(\Omega)$ be the solution of (23), and let $u_{h}\in\mathbb{V}_{h}$ be its finite element approximation, which solves (35). If $\nabla\mathbf{V}(\nabla u)\in L^{2}(\omega,\Omega)$, then
@@
where the hidden constant depends on $\sigma$, $\omega]_{A_{i(\Phi)}}$, and the characteristics of $\Phi$.
Proof 4.10.
The error bound (44) follows from the quasi-best approximation property (37) in conjunction with the error estimate (42).
@@
Note that Theorem 4.9 assumes $\nabla{\mathbf{V}}(\nabla u)\in L^{2}(\omega,\Omega)$. In the unweighted case, i.e., $\omega\equiv 1$, such results are proved in [1, 7, 8, 15, 18].
@@ 5 The obstacle problem
We now consider the constrained minimization of $\mathcal{J}$ defined in (1). In the same setting and with the same assumptions as in Section 3, we additionally assume that we are given $\psi\in C^{2}(\bar{\Omega})$ with $\psi\leq 0$ on $\partial\Omega$ and define
(45) $\mathcal{K}_{\psi}\coloneqq\left\{w\in W^{1,\Phi}_{0}(\omega,\Omega):w\geq\psi~a.e.~\emph{in }\Omega\right\}.$
We recall that $\Phi$ is an $N$-function that satisfies Assumption 3, and that the weight $\omega$ belongs to $A_{\Phi}$. The problem we are interested in is to find $u\in\mathcal{K}_{\psi}$ such that
$\mathcal{J}(u)\leq\mathcal{J}(v)\quad\forall v\in\mathcal{K}_{\psi}.$
A convex minimization argument [21, 26] shows that the minimizer can be equivalently characterized as the solution of the variational inequality: Find $u\in\mathcal{K}_{\psi}$ such that
@@
Since it will be useful for what follows, we define $\lambda\in\left(W^{1,\Phi}(\omega,\Omega)\right)^{\prime}$ as
@@
We note that (46) can thus be rewritten as follows: Find $u\in\mathcal{K}_{\psi}$ such that
@@
  • For every $\phi\in C_{0}^{\infty}(\Omega)$ that is nonnegative, i.e., $\phi\geq 0$ in $\Omega$, we have $\langle\lambda,\phi\rangle\geq 0.$
  • The following complementarity condition holds: $\langle\lambda,u-\psi\rangle=0.$
Proof 5.2.
Let $0\leq\phi\in C_{0}^{\infty}(\Omega)$. Define $v=u+\phi\in W^{1,\Phi}_{0}(\omega,\Omega)$. Since $v$ satisfies $v\geq u\geq\psi$ in $\Omega$, $v$ is admissible in (46). We can therefore replace $v$ in the variational inequality (46) and obtain $0\geq\langle\lambda,u-(u+\phi)\rangle=-\langle\lambda,\phi\rangle.$ This proves the first property. We also note that this shows that $\lambda$ defines a nonnegative distribution. The Riesz-Schwartz theorem (see [44, Chapter 1, §4, Théorème V] and [47, Chapter 1, Sec. 1.7, Theorem II]) then shows that $\lambda$ is a nonnegative Radon measure.
@@
Now let $\mathcal{N}\subset\Omega^{+}$ be open, and let $\phi\in C_{0}^{\infty}(\mathcal{N})$. Consider $v=u-\epsilon\phi\in W^{1,\Phi}_{0}(\omega,\Omega)$. For $\epsilon>0$, but sufficiently small, we have that $v\in\mathcal{K}_{\psi}$. In fact, we have that $u(x)>\psi(x)$ up to a null set in $\Omega^{+}$. Then, for a sufficiently small $\epsilon>0$, we obtain that $u(x)-\epsilon\phi(x)\geq\psi(x)$ in $\Omega$. If we replace this function $v$ in (46), we obtain
@@
[…]
@@ 5.1 Finite element discretization
To approximate the solution of (46), we retain the notation and constructions from Section 4. In addition, we need to introduce the positivity preserving interpolant $\wp_{h}:L^{1}(\Omega)\to\mathbb{V}_{h}$, which was originally developed in [12]. Let $\mathscr{V}=\{{\mathtt{v}}\}$ be the collection of vertices of $\mathscr{T}_{h}$ and let $\mathscr{V}_{\mathrm{in}}\coloneqq\mathscr{V}\cap\Omega$. Since $\mathbb{V}_{h}$ consists of piecewise linears, there is a bijection between $\mathscr{V}_{\mathrm{in}}$ and the canonical Lagrange basis $\{\phi_{\mathtt{v}}\}_{{\mathtt{v}}\in\mathscr{V}_{\mathrm{in}}}$ of $\mathbb{V}_{h}$. Given ${\mathtt{v}}\in\mathscr{V}_{\mathrm{in}}$, we let
@@
Standard arguments can be used to prove the existence and uniqueness of a solution of (51).
@@ 5.2 A priori error bounds
Let $u\in\mathcal{K}_{\psi}$ and $u_{h}\in\mathcal{K}_{h,\psi}$ be the solutions of (46) and (51), respectively. If $i(\Phi)>2$ and $\lambda\in L^{2}(\omega^{-1},\Omega)$, then we have
@@
Proof 5.4.
We first note that since $i(\Phi)>2$, Proposition 2.12 guarantees that $u\in W^{1,2}_{0}(\omega,\Omega)$. Next, we set some notation. We define the error $e:=u-u_{h}$ and the auxiliary function $z:=u-\wp_{h}u$. We note that $\wp_{h}u\in\mathcal{K}_{h,\psi}$.
@@
$\|{\mathbf{V}}(\nabla u)-{\mathbf{V}}(\nabla u_{h})\|_{L^{2}(\omega,\Omega)}^{2}\lesssim\int_{\Omega}\omega(x)\left[{\mathbf{A}}(\nabla u(x))-{\mathbf{A}}(\nabla u_{h}(x))\right]\cdot\nabla e(x)\,\mbox{\rm d}x\eqqcolon\mathfrak{D}.$
We further decompose $\mathfrak{D}$ as follows:
@@
    $\leq\|\lambda\|_{L^{2}(\omega^{-1},\Omega)}\|\left[u-\psi\right]-\wp_{h}\left[u-\psi\right]\|_{L^{2}(\omega,\Omega)}.$
@@
$\mathfrak{L}\leq C\delta\|{\mathbf{V}}(\nabla u)-{\mathbf{V}}(\nabla u_{h})\|_{L^{2}(\omega,\Omega)}^{2}+C_{\delta}\int_{\Omega}\omega(x)\Phi_{|\nabla u(x)|}(|\nabla z(x)|)\,\mbox{\rm d}x.$
If we insert the estimates for $\mathfrak{L}$ and $\mathfrak{N}$ into the bound in which $\mathfrak{D}$ is defined, we obtain
$\|{\mathbf{V}}(\nabla u)-{\mathbf{V}}(\nabla u_{h})\|_{L^{2}(\omega,\Omega)}^{2}\leq C\delta\|{\mathbf{V}}(\nabla u)-{\mathbf{V}}(\nabla u_{h})\|_{L^{2}(\omega,\Omega)}^{2}\\ +C_{\delta}\int_{\Omega}\omega(x)\Phi_{|\nabla u(x)|}(|\nabla z(x)|)\,\mbox{\rm d}x+C\|\lambda\|_{L^{2}(\omega^{-1},\Omega)}\|\left[u-\psi\right]-\wp_{h}\left[u-\psi\right]\|_{L^{2}(\omega,\Omega)}.$
The term $\|{\mathbf{V}}(\nabla u)-{\mathbf{V}}(\nabla u_{h})\|_{L^{2}(\omega,\Omega)}^{2}$ appearing on the right hand side of the previous bound can be absorbed on the left hand side if $\delta$ is chosen carefully. This shows that
$\|{\mathbf{V}}(\nabla u)-{\mathbf{V}}(\nabla u_{h})\|_{L^{2}(\omega,\Omega)}^{2}\lesssim\|\lambda\|_{L^{2}(\omega^{-1},\Omega)}\|\left[u-\psi\right]-\wp_{h}\left[u-\psi\right]\|_{L^{2}(\omega,\Omega)}\\ +\int_{\Omega}\omega(x)\Phi_{|\nabla u(x)|}(|\nabla z(x)|)\,\mbox{\rm d}x,\qquad z=u-\wp_{h}u.$
Use the equivalences (30)–(31) one last time to obtain the assertion and conclude.
@@ 5.3 The positivity preserving interpolant
Note that up to this point no property other than positivity has been required of $\wp_{h}$. In the following, however, the stability and approximation properties of $\wp_{h}$ are required. Ideally, we would like to have an operator that is a projection and satisfies (38). Note, however, that $\wp_{h}$ is not a projection. In fact, it is not possible to construct a positivity preserving projection; see [41]. For this reason, we need to prove suitable properties for $\wp_{h}$.
@@
    $\wp_{h}$ is weighted Orlicz stable: Let $\Phi$ be an $N$-function that satisfies Assumption 3, and let $\omega\in A_{\Phi}$. Then, for every $T\in\mathscr{T}_{h}$ and for all $a\geq 0$,
@@
    $\wp_{h}$ has weighted approximation properties: If $\omega\in A_{2}$ and $w\in W^{1,2}_{0}(\omega,\Omega)$, then, for all $T\in\mathscr{T}_{h}$,
@@
Proof 5.6.
We examine each statement individually.
@@
    Since $w_{|\mathcal{B}_{\mathtt{v}}}\in\mathbb{P}_{1}$, we write $w(x)=a+\mathbf{b}\cdot(x-{\mathtt{v}})$ for $a\in\mathbb{R}$ and $\mathbf{b}\in\mathbb{R}^{d}$. We have
@@
    This follows from the repetition of the previous calculation for each vertex of $T$, which forms an unisolvent set for $\mathbb{P}_{1}(T)$.
@@
    A suitable choice of $\mathfrak{p}$ shall then imply the thesis.
    To prove (52), we proceed differently depending on whether $T$ is an interior element or $T$ touches the boundary. If $T$ is an interior element, i.e., $T\cap\partial\Omega=\emptyset$, then local invariance and stability in $L^{2}(\omega,T)$ imply the result. On the other hand, if the element $T$ touches the boundary we proceed as follows. Let $\{{\mathtt{v}}_{i}^{T}\}_{i=0}^{d}$ be the collection of vertices of $T$. Since $(\wp_{h}w-\mathfrak{p})_{|T}\in\mathbb{P}_{1}$, we can write
@@
    We now use that $\|\phi_{\mathtt{v}}\|_{L^{\infty}(\Omega)}=1$ for all ${\mathtt{v}}\in\mathscr{V}$ to obtain the estimate
@@
    On the other hand, if ${\mathtt{v}}_{i}^{T}\in\partial\Omega$, then $\wp_{h}w({\mathtt{v}}_{i}^{T})=0$ by definition. Let $F\subset T$ be the (d1)(d-1)-dimensional subsimplex containing ${\mathtt{v}}_{i}^{T}$ such that $F\subset\partial\Omega$. We thus have
@@
Proposition 5.7 (interpolation error estimate).
Let $\Phi$ be an $N$-function that satisfies Assumption 3, and let $\omega\in A_{i(\Phi)}(\Omega)$. Let us assume that $\mathscr{T}_{h}=\{T\}$ is such that no simplex $T$ has more than one (d1)(d-1)-dimensional subsimplex on $\partial\Omega$. If $w\in W^{1,\Phi}_{0}(\omega,\Omega)$ is such that $\nabla{\mathbf{V}}(\nabla w)\in L^{2}(\omega,\Omega)$ and $h$ is sufficiently small, then
@@
where the constant depends on $\omega]_{A_{i(\Phi)}}$ and $\|\omega\|_{L^{\infty}({\mathcal{G}})}$, but is independent of $h$.
Proof 5.8.
@@
Let $T\in\mathscr{T}_{h}^{\mathrm{in}}$. The local invariance of $\wp_{h}$, proved in Lemma 5.5, implies that $\wp_{h}\mathfrak{p}_{|T}=\mathfrak{p}_{|T}$. Using the equivalences (30)–(31) and the weighted Orlicz stability of $\wp_{h}$, which was derived in Lemma 5.5, we obtain the estimate
$\mathbf{I}\mathbf{I}$ $\lesssim\int_{T}\omega(x)\Phi_{|\nabla\mathfrak{p}|}(|\nabla\wp_{h}(w-\mathfrak{p})(x)|)\,\mbox{\rm d}x\lesssim\int_{S_{T}}\omega(x)\Phi_{|\nabla\mathfrak{p}|}(|\nabla(w-\mathfrak{p})(x)|\,\mbox{\rm d}x$
@@
It follows that for every $T\in\mathscr{T}_{h}^{\mathrm{in}}$ the following bound applies:
@@
where $\mathfrak{p}\in\mathbb{P}_{1}(\mathbb{R}^{d})$ is arbitrary. Let $\mathbf{Q}\in\mathbb{R}^{d}$ and choose $\mathbf{z}=\nabla\mathfrak{p}\in\mathbb{R}^{d}$, so that ${\mathbf{V}}(\mathbf{z})=\mathbf{Q}$. This is possible because $\mathbf{V}$} is surjective. This gives us for every $T\in\mathscr{T}_{h}^{\mathrm{in}}$
@@
We now consider $T\in\mathscr{T}_{h}^{\partial}$. If $h$ is sufficiently small, it is not a mistake to assume that $T\subset{\mathcal{G}}$. Recall that $\mathcal{G}$} is as in Definition 2.1. Let $F\subset T$ be the unique (d1)(d-1)-dimensional subsimplex such that $F=T\cap\partial\Omega$ and denote by $\boldsymbol{\zeta}$ the unit outer normal to $T$ on $F$. Now let $\{{\mathtt{v}}_{i}^{T}\}_{i=0}^{d}$ be the set of vertices of $T$ and assume that $\{{\mathtt{v}}_{i}^{T}\}_{i=0}^{d}$ is numbered such that $F=\operatorname{conv}\{{\mathtt{v}}_{i}^{T}\}_{i=0}^{d-1}$. Define $\mathfrak{p}\in\mathbb{P}_{1}(\mathbb{R}^{d})$ as $\mathfrak{p}(x)=b\boldsymbol{\zeta}\cdot(x-{\mathtt{v}}_{0}^{T})$, where $b\in\mathbb{R}$ is to be chosen. We note that, by construction, $\mathfrak{p}_{|F}=0$. Moreover, we have by definition that $\wp_{h}\mathfrak{p}_{|F}=0$. Finally, the symmetry of $\wp_{h}$ proved in Lemma 5.5 implies that $\wp_{h}\mathfrak{p}({\mathtt{v}}_{d}^{T})=\mathfrak{p}({\mathtt{v}}_{d}^{T})$. From this follows that $\wp_{h}\mathfrak{p}_{|T}=\mathfrak{p}_{|T}$.
Based on the constructions described above, the weighted Orlicz stability of $\wp_{h}$ (see Lemma 5.5) allows us to treat $\mathbf{I}\mathbf{I}$ as before and obtain that, for every $T\in\mathscr{T}_{h}^{\partial}$,